# ”Security for Everyone” in Finite Blocklength IRS-aided Systems With Perfect and Imperfect CSI

### v2

@@ Abstract
Provisioning secrecy for all users, given the heterogeneity in their channel conditions, locations, and the unknown location of the attacker/eavesdropper, is challenging and not always feasible. The problem is even more difficult under finite blocklength constraints that are popular in ultra-reliable low-latency communication (URLLC) and massive machine-type communications (mMTC). This work takes the first step to guarantee secrecy for all URLLC/mMTC users in the finite blocklength regime (FBR) where intelligent reflecting surfaces (IRS) are used to enhance legitimate users’ reception and thwart the potential eavesdropper (Eve) from intercepting. To that end, we aim to maximize the minimum secrecy rate (SR) among all users by jointly optimizing the transmitter’s beamforming and IRS’s passive reflective elements (PREs) under the FBR latency constraints. The resulting optimization problem is non-convex and even more complicated under imperfect channel state information (CSI). To tackle it, we linearize the objective function, and decompose the problem into sequential subproblems. When perfect CSI is not available, we use the successive convex approximation (SCA) approach to transform imperfect CSI related semi-infinite constraints into finite linear matrix inequalities (LMI). We prove that our proposed algorithm’s converges to a locally optimal solution with low computational complexity thanks to our closed-form linearization approach. This makes the solution scalable for large IRS deployments. Extensive simulations with practical settings show that our approach can ensure secure communication for all users while satisfying FBR constraints even with only imperfect CSI.
@@
Finite blocklength regime, ultra reliable and low latency communication, intelligent reflective surfaces (IRS), fairness max-min optimization, and secrecy rate (SR).
@@ I Introduction
Intelligent reflective surface (IRS) has emerged as a promising technology, garnering significant interest due to its ability to enhance/tailor the passive radio environment. IRSs consist of passive reflective elements (PREs) equipped with phase shift controllers, allowing them to intelligently manipulate reflected signals [1]. By optimizing signal reflections, IRSs improve reception by creating favorably reflected multi-path signals at receivers. Thanks to their cost-effective design and convenient deployment typically on the facades of high-rise buildings, IRSs hold immense potential for various applications, particularly in urban areas where line-of-sight channels between transmitters (Tx) and receivers (Rx) frequently face obstructions [2]. Particularly, for the Internet of Things (IoT) devices that are limited computing capability and battery/energy, the IRS has gained paramount attention aiming to enhance both spectral and energy efficiency [3].
@@ I-A Related Works and Motivations
Another potential application of IRSs is to enhance the security/privacy of users by purposely manipulating reflected signals from the Tx to facilitate the signal reception at legitimate users while maximizing the multi-user interference/degrading the signals at potential eavesdroppers. In [4], the authors studied the design of a single user IRS-aided system to maximize the legitimate user’s secrecy rate (SR), which is defined as the difference between the rate of the legitimate channel and that of the channel from the transmitter to a potential eavesdropper. The authors maximize the single-user SR by jointly optimizing the beamformer at the transmitter and the IRS. A multi-input-single-output (MISO) IRS-aided system with the presence of multiple eavesdroppers was considered in [5]. The transmitter introduced artificial noise (AN) as a countermeasure to enhance user security. In [6], the authors studied a cell-free IRS-aided network, where they aimed to enhance the user’s secrecy by maximizing the weighted sum of the users’ secrecy rate (WSSR) by jointly optimizing the beamforming vector and the IRS’s PREs. The alternating optimization (AO) algorithm was used to decouple the beamforming factor and the IRS’s PREs, and the resulting problem was tackled using semi-definite relaxation (SDR) and continuous convex approximation. Shi et al. in [7] investigated the SR in a multi-input multi-output (MIMO) IRS-aided system with the presence of a single eavesdropper. When the channel state information (CSI) from the IRS to the users/receivers is unknown or imperfect, the authors in [8] studied an IRS-aided multi-user system, where they formulated a secrecy sum rate (SSR) maximization problem with the eavesdropper’s channel partially known to the receiver. It is worth mentioning that the SSR problem does not guarantee the secrecy for all users.
@@
In parallel, low latency communication (URLLC) and massive machine type communication (mMTC), both relying on finite blocklength (short packet) regime/communication (FBR), have been envisioned as key technologies to serve critical missions of IoT. Since FBR often requires a stricter design approach than the long blocklength regime/communication (LBR) systems[9], maintaining high-reliability communication is more challenging due to the lower channel coding gain. Moreover, in mMTC and URLLC applications such as intelligent transportation, leakage information could expose the user’s location or identity. As a potential solution to secure URLLC and mMTC, the use of IRS has recently attracted paramount interest. IRS links exhibit low latency that is beneficial for the FBR strict requirements [10]. Zhao et al. studied the information freshness in a single user FBR-IRS-aided system with the presence of an eavesdropper in [10]. The authors derived a closed-form expression for the upper bound secrecy outage probability under statistical CSI. The single user’s secrecy outage is minimized by jointly optimizing the blocklength and the IRS PREs. In [11], the authors investigated the problem of transmission power minimization in an IRS-aided system under the FBR regime with the presence of multiple eavesdroppers, by jointly optimizing the beamforming vector, the AN, and the IRS PREs. In [12], the authors derived the outage probability and block error rate (BLER) from users and eavesdroppers. In [13], the authors studied the SSR maximization problem for an FBR-IRS-aided system by jointly optimizing the beamforming vector, the IRS PREs, and the blocklength. The optimization problem was tackled using Riemannian conjugate gradient (RCG) based algorithm. However, the assumption that Alice/Tx can obtain Eve’s CSI, as stated before, is not practical. Note that most earlier work focused only on single-user systems. The SSR maximization approach in LBR-IRS systems is neither applicable to nor can guarantee the secrecy for all users in FBR-IRS systems due to the additional latency requirements of the FBR.
@@
Motivated by the above, this work takes the first step to guarantee secrecy for all URLLC/mMTC users in FBR-IRS-aided systems. Specifically, we consider a popular case where IRSs are used to enhance/aid the signal reception at legitimate users (from a transmitter) under the presence of a potential eavesdropper while the direct/line-of-sight channel between the transmitter to all users and eavesdropper does not exist (e.g., due to blockages). [^1] In reality, it is impractical to consider the underlying problem if the eavesdropper has a line-of-sight channel with the transmitter while the users do not. In such cases, a friendly jammer should be used to protect users’ privacy [14]. We then maximize the minimum SR among all the users while maintaining the FBR constraints by optimizing the transmit beamforming vector and the IRS’s PREs. We consider both cases with perfect and imperfect CSI from the IRS to the users and the eavesdropper. For the perfect CSI, we assume that the transmitter can obtain the correct CSI of all users and the eavesdropper.This is the case when one of the legitimate users gets compromised and acts as an eavesdropper, trying to eavesdrop on other users, hence its CSI would be available to Alice.
@@
When only partial or erroneous CSI (from the IRS to the users and the eavesdropper) is available, the problem is even more challenging due to the semi-infinite constraints invoked by the imperfect CSI and the non-convexity nature of the dispersion factors imposed by the FBR transmission constraints. To tackle the problem, we first introduce slack variables to deal with the transmit beamforming vectors and the IRS’s PREs coupling within the objective function. Secondly, we leverage the successive convex approximation (SCA) technique [15], and the $\mathcal{S}$-procedure [16] to convert the semi-infinite constraints into linear matrix inequalities (LMI). Then, we propose a penalty convex-concave procedure (PCCP) [17] to tackle the unit modulus constraint (UMC) of the IRS. Extensive simulations with practical settings show that maximizing the minimum SR among all the users can achieve a better chance of ensuring secure communications for all users (subject to the location of the eavesdropper) under the FBR constraints even under imperfect CSI conditions.
@@ I-B Contributions
  • We take the first step to guarantee secrecy for all URLLC/mMTC users in the finite blocklength regime (FBR) where IRSs are used to enhance legitimate users’ reception while thwarting the potential Eve from intercepting. To that end, we aim to maximize the minimum SR among all users by jointly optimizing the transmitter’s beamforming and IRS’s passive reflective elements (PREs) while meeting FBR latency constraints. The resulting optimization problem is non-convex and even more complicated under imperfect channel state information (CSI).
  • When the CSI is available, we tackle the above non-convex problem by linearizing its objective function with tractable approximation functions, leading to a computationally efficient algorithm. To tackle UMC, unlike traditional methods, i.e. conventional SDR, we directly optimize the PREs arguments to provide a low-complexity solution that can scale with large IRSs. The resulting solution is proved to converge to a locally optimal solution of the original non-convex problem.
  • When perfect CSI is not available, we use the successive convex approximation (SCA) approach to transform imperfect CSI related semi-infinite constraints into finite linear matrix inequalities (LMI). We prove that our proposed algorithm converges to a locally optimal solution with low computational complexity thanks to our closed-form linearization approach. This makes the solution scalable for large IRS deployments.
  • Eventually, we perform extensive simulations with practical settings. The simulations results show that our approach can ensure secure communication for all users while satisfying FBR constraints even under imperfect CSI. Additionally, the impact of the reflected channel’s imperfect CSI on the system is also evaluated.
@@
This paper uses the following notation: bold letters denote vectors and matrices. $\boldsymbol{I}_{M}$ denotes an M dimensional identity matrix. Diag($n_{1},\dots,n_{n}$) denotes the diagonal matrix with diagonal entries of $\{n_{1},\dots,n_{n}\}$. The symbols $\Re$ and $\mathbb{C}$ represent the real and the complex field, respectively. ${\cal C}(0,\bar{z})$ denotes the circular Gaussian random variable with zero mean and variance $\bar{z}$. For matrices $\mathbf{C}$ and $\mathbf{D}$, $\langle\mathbf{C},\mathbf{D}\rangle\triangleq\mbox{trace}\left(\mathbf{C}^{H}\mathbf{D}\right)$. For matrix $\mathbf{A}$, $\Re\left\{\mathbf{A}\right\}$ denotes the real part, $\langle\mathbf{A}\rangle\triangleq\mbox{trace}(\mathbf{A})$, the symbol $\|\mathbf{A}\|_{1}$ denotes the 1-norm, $\|\mathbf{A}\|$ denotes the Frobenius norm, $\mathbf{A}^{*}$ denotes the conjugate, $\mathbf{A}^{H}$ denotes the Hermitian (conjugate transpose), $\lambda_{\text{max}}\left(\mathbf{A}\right)$ denotes the maximal eigenvalue, $\angle(\mathbf{A})$ denotes its argument, and $\mathbf{A}\succeq 0$ means positive semi-definite.
@@ II System Model
$y_{i}\triangleq\boldsymbol{h}_{i}(\boldsymbol{\theta}){\sum}_{k=1}^{\mathcal{K}}\boldsymbol{w}_{k}s_{k}+n_{i},i\in\{k,e\},$ (1)
@@
where $\boldsymbol{h}_{i}(\boldsymbol{\theta})\in\mathbb{C}^{1\times M}$ is the cascaded channel gain from Alice to $i\in\{k,e\}$, $\boldsymbol{w}_{k}\in\mathbb{C}^{M\times 1}$ is the beamforming vector applied for user-$k$, $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{N})^{T}\in[0,2\pi)^{N}$ denotes the phase shift vector the IRS’s PREs, $n_{i}$ is the zero-mean Additive White Gaussian Noise (AWGN) with power density $\sigma_{i}$ for $i\in\{k,e\}$. The cascade channel gain $\boldsymbol{h}_{i}(\boldsymbol{\theta})$ from Alice to $i\in\{k,e\}$, can be written in terms of the channel gain from Alice to the IRS and the channel gain from the IRS to the user or Eve as follows:
$\boldsymbol{h}_{i}(\boldsymbol{\theta})\triangleq{\boldsymbol{u}}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\triangleq{\boldsymbol{u}}_{i}\sum_{n=1}^{N}\exp(j\theta_{n})\Xi_{n}\boldsymbol{G}_{{}_{\text{AR}}},$ (2)
@@
where $\boldsymbol{\Xi}_{n}$ is an $N\times N$ matrix with all zeros except for its $(n,n)$ entry which is 1, and $\boldsymbol{\Phi}=\text{Diag}(e^{j\boldsymbol{\theta}})$.
@@
The CSI from Alice to the IRS can be estimated by calculating the angle of arrival and departure [20]. However, the reflected channel’s CSI from the IRS to the IoT users is much more challenging to obtain due to the passive nature of the IRS and the mobility/varying nature of the users’ environment and location [21]. For Eve, it is also hardly possible to pinpoint its location or its accurate CSI from the IRS. To account for the CSI imperfection, we adopt the bounded CSI model in which the reflected channel from the IRS to the users and Eve can be expressed as [22]:
$\boldsymbol{u}_{i}$ $\triangleq\widehat{\boldsymbol{u}}_{i}+\Delta\boldsymbol{u}_{i},\forall i\in\{k,e\},$ (3a)
$\omega_{i}$$\triangleq\{\|\Delta\boldsymbol{u}_{i}\|_{2}\leq\Omega_{i}\},\forall i\in\{k,e\},$ (3b)
@@
where $\widehat{\boldsymbol{u}}_{i}$ denotes the (imperfect) estimated channel vector, $\Delta\boldsymbol{u}_{i}$ represents the channel estimation error of the corresponding estimation, $\omega_{i}$ is a set for all possible channel estimation errors, and $\Omega_{i}$ is the radii of the uncertainty regions as known to Alice. Hence, (2) can be reformulated as:
@@
$\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\triangleq\left(\widehat{{\boldsymbol{u}}}_{i}+\Delta\boldsymbol{u}_{i}\right)\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}.$ (4)
@@
In the sequel, we deal with perfect and imperfect CSI from the IRS to the users and Eve. For the former, we assume that the CSI of all the users and Eve can be accurately obtained by Alice [23, 24]. As aforementioned, in this case we assume that Eve acts as a legitimate user in one period/slot but then tries to eavesdrop on other users. Using different well-established channel estimation methods, such as the anchor-assisted channel estimation approach [25], Alice can practically obtain the CSI of all the users. In this case, the reflected channel from the IRS to the users and Eve can be captured by setting $\Delta\boldsymbol{u}_{i}$ to zero in (3a), i.e., $\boldsymbol{u}_{i}=\widehat{\boldsymbol{u}}_{i},i\in\{k,e\}$. In the second case, only partial/imperfect CSI $\widehat{\boldsymbol{u}}_{i}$ is available. In this case, one can vary the magnitude of $\Delta\boldsymbol{u}_{i}$, i.e., the radii $\Omega_{i}$ of the uncertainty region to capture different levels of CSI imperfection.
@@
$\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{|\boldsymbol{h}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{k}|^{2}}{\rho_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (5)
$\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{|\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{k}|^{2}}{\widehat{\rho}_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (6)
where $\rho_{i}\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}|\boldsymbol{h}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i}$, and $\widehat{\rho}_{i}\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}|\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i}$. Under the presence of Eve, the closed-form expressions for the FBR-SR of the user-$k$ under perfect and imperfect IRS to users/Eve CSI are defined as, respectively, [26]:
$\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\left[C_{k}-\xi_{k}\sqrt{V_{k}}-\left(C_{e}+\xi_{e}\sqrt{V_{e}}\right)\right]^{+},$ (7)
$\widehat{\mathcal{S}}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\left[\widehat{C}_{k}-\xi_{k}\sqrt{\widehat{V}_{k}}-\left(\widehat{C}_{e}+\xi_{e}\sqrt{\widehat{V}_{e}}\right)\right]^{+},$ (8)
where $[x]^{+}\triangleq\text{max}[0,x]$, while the user’s data rate, and the eavesdropping rate of Eve decoding $s_{k}$, are respectively given by,

### v3

@@ Abstract
Provisioning secrecy for all users, given the heterogeneity in their channel conditions, locations, and the unknown location of the attacker/eavesdropper, is challenging and not always feasible. The problem is even more difficult under finite blocklength constraints that are popular in ultra-reliable low-latency communication (URLLC). This work takes the first step to guarantee secrecy for all URLLC users in the finite blocklength regime (FBR) where intelligent reflecting surfaces (IRS) are used to enhance legitimate users’ reception and thwart the potential eavesdropper (Eve) from intercepting. To this end, we aim to maximize the minimum secrecy rate (SR) across all users by jointly optimizing the transmitter’s beamforming vectors and the IRS’s passive reflective elements (PREs) under FBR latency constraints. In the FBR-SR expression, the square-root dispersion factor under the FBR, couples with the latency and reliability constraints, making the problem significantly more challenging than the same problem in the long blocklength regime (LBR). Under imperfect CSI, the problem becomes even more challenging due to nonconvex, semi-infinite constraints arising from channel estimation errors. To address this, we approximately linearize the objective function, constraints and decompose the problem into sequential subproblems. We then apply the successive convex approximation (SCA), and the $\mathcal{S}$-procedure approach to transform imperfect CSI-related semi-infinite constraints into finite linear matrix inequalities (LMIs) that can be efficiently solved. We prove that our proposed algorithms converge to a locally optimal solution with low computational complexity thanks to our closed-form linearization approach. This makes the solution scalable for large IRS deployments. Extensive simulations with practical settings show that our approach can ensure secure communication for all users while satisfying FBR constraints even with only imperfect CSI.
@@
Finite blocklength regime, ultra-reliable and low-latency communication, intelligent reflective surfaces (IRS), fairness max-min optimization, and secrecy rate (SR).
@@ I Introduction
Intelligent reflective surface (IRS) has emerged as a promising technology, garnering significant interest due to its ability to enhance/tailor the passive radio environment. IRSs consist of passive reflective elements (PREs) equipped with phase shift controllers, allowing them to intelligently manipulate reflected signals [1]. By optimizing signal reflections, IRSs improve reception by creating favorably reflected multi-path signals at the receivers. Thanks to their cost-effective design and convenient deployment typically on the facades of high-rise buildings, IRSs hold immense potential for various applications, particularly in urban areas where line-of-sight channels between transmitters (Tx) and receivers (Rx) frequently face obstructions [2]. Particularly, for the Internet of Things (IoT) devices that have limited computing capability and battery/energy, the IRS has gained paramount attention aiming to enhance both spectral and energy efficiency [3].
@@ I-A Related Works and Motivations
Another potential application of IRSs is to enhance the security/privacy of users by purposely manipulating reflected signals from the Tx to facilitate the signal reception at legitimate users while maximizing the multi-user interference/degrading the signals at potential eavesdroppers. In [4], the authors studied the design of a single user IRS-aided system to maximize the legitimate user’s secrecy rate (SR), which is defined as the difference between the rate of the legitimate channel and that of the channel from the transmitter to a potential eavesdropper. The authors maximize the single-user SR by jointly optimizing the beamformer at the transmitter and the IRS. When the channel state information (CSI) from the IRS to the users/receivers is unknown or imperfect, the authors in [5] studied an IRS-aided multi-user system, where they formulated a secrecy sum rate (SSR) maximization problem with the eavesdropper’s channel partially known to the receiver. It is worth mentioning that the SSR problem does not guarantee the secrecy for all users.
@@
In parallel, low-latency communication (URLLC), which relies on the finite blocklength (short packet) regime of communication (FBR), has been envisioned as a key technology to serve the critical missions of IoT. Since FBR often requires a stricter design approach than the long blocklength regime/communication (LBR) systems[6], maintaining high-reliability communication is more challenging due to the lower channel coding gain. Moreover, in URLLC applications such as intelligent transportation, leakage information could expose the user’s location or identity. As a potential solution to secure URLLC, the use of IRS has recently attracted paramount interest. IRS links exhibit low latency that is beneficial for the FBR strict requirements [7]. Zhao et al. studied the information freshness in a single-user FBR-IRS-aided system with the presence of an eavesdropper in [7]. The authors derived a closed-form expression for the upper bound secrecy outage probability under statistical CSI. The single user’s secrecy outage is minimized by jointly optimizing the blocklength and the IRS PREs. The authors in [8] studied the secrecy performance of an IRS-assisted URLLC system, where an FBR coding scheme was proposed to secure the link, showing that secrecy is achievable beyond a certain blocklength threshold. However, it does not consider the joint optimization of secrecy rate and transmission parameters, which is crucial for practical deployment. Nevertheless, the assumption that Alice/Tx can obtain Eve’s CSI, as stated before, is not always feasible. Note that most earlier work focused only on single-user systems. The SSR maximization approach in LBR-IRS systems is neither applicable to nor can it guarantee the secrecy for all users in FBR-IRS systems due to the additional latency requirements of the FBR.
@@
Guaranteeing secrecy in FBR is significantly more challenging than in conventional LBR applications. Unlike LBR, where sufficiently long codewords can asymptotically drive both error probability and leakage to zero, these factors remain strictly non-zero in FBR due to the coding limits imposed by short blocklengths [9]. In particular, the square-root of the channel dispersion in the FBR-SR expression directly captures the effect of finite blocklength on the decoding error probability, making the SR a non-linear function. Consequently, the optimization problem is significantly more complex, as these non-zero error and leakage terms introduce additional nonconvex penalties that must be jointly considered with power and beamforming constraints. Motivated by the above, this work takes the first step to guarantee secrecy for all URLLC users in FBR-IRS-aided systems. Unlike existing work, which only considers secrecy enhancement in LBR-IRS-aided systems, or FBR secrecy improvement with full CSI [10], we tackle the problem of secrecy guarantee in FBR-IRS-aided systems under perfect/imperfect IRS to users’/Eve’s CSI. To this end, we consider a popular case where IRSs are used to enhance/aid the signal reception at legitimate users (from a transmitter) under the presence of a potential eavesdropper, while the direct/line-of-sight channel between the transmitter to all users and eavesdropper does not exist (e.g., due to blockages). [^1] In reality, it is impractical to consider the underlying problem if the eavesdropper has a line-of-sight channel with the transmitter while the users do not. In such cases, a friendly jammer should be used to protect users’ privacy [11]. We then maximize the minimum SR among all the users given the FBR constraints by optimizing the transmit beamforming vector and the IRS’s PREs. We consider both cases with perfect and imperfect CSI from the IRS to the users and the eavesdropper. For the perfect CSI, we assume that the transmitter can obtain the correct CSI of all users and the eavesdropper. This is the case when one of the legitimate users gets compromised and acts as an eavesdropper, trying to eavesdrop on other users; hence, its CSI would be available to Alice.
@@
When only partial or erroneous CSI (from the IRS to the users and Eve) is available, the problem becomes even more challenging. In particular, the semi-infinite constraints introduced by imperfect CSI are embedded within the dispersion term of the FBR-SR expression, which involves both the inverse Q-function and the square-root of the channel dispersion. This embedding makes the decoding error probability highly sensitive to channel estimation errors [12], while the square-root further increases the nonlinearity. To tackle the problem, we first introduce slack variables to deal with the transmit beamforming vectors and the IRS’s PREs coupling within the objective function. Secondly, we leverage the successive convex approximation (SCA) technique [13], and the $\mathcal{S}$-procedure [14] to convert the semi-infinite constraints into linear matrix inequalities (LMI). Although the SCA framework with the $\mathcal{S}$-procedure converts the semi-infinite constraints into LMIs, they remain nonconvex. We address this by applying a first-order Taylor approximation to linearize the nonconvex components, yielding a convex reformulation solvable at each iteration. Then, we propose a penalty convex-concave procedure (PCCP) [15] to tackle the unit modulus constraint (UMC) of the IRS. Extensive simulations with practical settings show that maximizing the minimum SR among all the users can achieve a better chance of ensuring secure communications for all users (subject to the location of the eavesdropper) under the FBR constraints even under imperfect CSI conditions.
@@ I-B Contributions
  • We take the first step to guarantee secrecy for all URLLC users in the finite blocklength regime (FBR) where IRSs are used to enhance legitimate users’ reception while thwarting the potential Eve from intercepting. To that end, we aim to maximize the minimum SR among all users by jointly optimizing the transmitter’s beamforming and the IRS’s passive reflective elements (PREs), while meeting the FBR latency constraints. The resulting optimization problem is nonconvex and even more complicated under imperfect channel state information (CSI).
  • When the CSI is available, we tackle the above nonconvex problem by linearizing its objective function with tractable approximation functions, leading to a computationally efficient algorithm. To tackle UMC, unlike traditional methods, i.e., conventional SDR, we directly optimize the PREs arguments to provide a low-complexity solution that can scale with large IRSs. The resulting solution is proved to converge to a locally optimal solution of the original nonconvex problem.
  • When perfect CSI is not available, we use the SCA approach to transform imperfect CSI-related semi-infinite constraints into finite LMI. We prove that our proposed algorithm converges to a locally optimal solution with low computational complexity thanks to our closed-form linearization approach. This makes the solution scalable for large IRS deployments.
  • Eventually, we perform extensive simulations with practical settings. The simulation results show that our approach can ensure secure communication for all users while satisfying FBR constraints even under imperfect CSI. Additionally, the impact of the reflected channel’s imperfect CSI on the system is also evaluated.
@@
This paper uses the following notation: bold letters denote vectors and matrices. $\boldsymbol{I}_{M}$ denotes an M dimensional identity matrix. Diag($m_{1},\dots,m_{m}$) denotes the diagonal matrix with diagonal entries of $\{m_{1},\dots,m_{m}\}$. The symbols $\Re$ and $\mathbb{C}$ represent the real and the complex field, respectively. ${\cal C}(0,\bar{z})$ denotes the circular Gaussian random variable with zero mean and variance $\bar{z}$. For matrices $\mathbf{C}$ and $\mathbf{D}$, $\langle\mathbf{C},\mathbf{D}\rangle\triangleq\mbox{trace}\left(\mathbf{C}^{H}\mathbf{D}\right)$. For matrix $\mathbf{A}$, $\Re\left\{\mathbf{A}\right\}$ denotes the real part, $\langle\mathbf{A}\rangle\triangleq\mbox{trace}(\mathbf{A})$, the symbol $\|\mathbf{A}\|_{1}$ denotes the 1-norm, $\|\mathbf{A}\|$ denotes the Frobenius norm, $\mathbf{A}^{*}$ denotes the conjugate, $\mathbf{A}^{H}$ denotes the Hermitian (conjugate transpose), $\lambda_{\text{max}}\left(\mathbf{A}\right)$ denotes the maximal eigenvalue, $\angle(\mathbf{A})$ denotes its argument, and $\mathbf{A}\succeq 0$ means positive semi-definite.
@@ II System Model
$g_{i}\triangleq\boldsymbol{h}_{i}(\boldsymbol{\theta}){\sum}_{k=1}^{\mathcal{K}}\boldsymbol{w}_{k}s_{k}+\bar{n}_{i},i\in\{k,e\},$ (1)
@@
where $\boldsymbol{h}_{i}(\boldsymbol{\theta})\in\mathbb{C}^{1\times N}$ is the cascaded channel gain from Alice to $i\in\{k,e\}$, $\boldsymbol{w}_{k}\in\mathbb{C}^{N\times 1}$ is the beamforming vector applied for user-$k$, $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{m})^{T}\in[0,2\pi)^{M}$ denotes the phase shift vector the IRS’s PREs, $\bar{n}_{i}$ is the zero-mean Additive White Gaussian Noise (AWGN) with power density $\sigma_{i}$ for $i\in\{k,e\}$. The cascade channel gain $\boldsymbol{h}_{i}(\boldsymbol{\theta})$ from Alice to $i\in\{k,e\}$, can be written in terms of the channel gain from Alice to the IRS and the channel gain from the IRS to the user or Eve as follows:
$\boldsymbol{h}_{i}(\boldsymbol{\theta})\triangleq{\boldsymbol{l}}_{i}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}}\triangleq{\boldsymbol{l}}_{i}\sum_{m=1}^{M}\exp(j\theta_{m})\boldsymbol{\Gamma}_{m}\boldsymbol{L}_{\text{AR}},$ (2)
@@
where $\boldsymbol{\Gamma}_{m}$ is an $M\times M$ matrix with all zeros except for its $(n,n)$ entry which is 1, and $\boldsymbol{\Phi}=\text{Diag}(e^{j\boldsymbol{\theta}})$.
@@
Given that the IRS is deployed at a known location, such as the facade of a high-rise building, the CSI between Alice and the IRS can be accurately estimated by exploiting the angles of arrival and departure [18], or through advanced channel estimation techniques [19, 20]. However, the reflected channel’s CSI from the IRS to the IoT users is much more challenging to obtain due to the passive nature of the IRS and the mobility/varying nature of the users’ environment and location [21]. For Eve, it is also hardly possible to pinpoint its location or its accurate CSI from the IRS. To account for the CSI imperfection, we adopt the bounded CSI model in which the reflected channel from the IRS to the users and Eve can be expressed as [22]:
$\boldsymbol{l}_{i}$ $\triangleq\widehat{\boldsymbol{l}}_{i}+\Delta\boldsymbol{l}_{i},\forall i\in\{k,e\},$ (3a)
$\omega_{i}$ $\triangleq\{\|\Delta\boldsymbol{l}_{i}\|_{2}\leq\Omega_{i}\},\forall i\in\{k,e\},$ (3b)
@@
where $\widehat{\boldsymbol{l}}_{i}$ denotes the (imperfect) estimated channel vector, $\Delta\boldsymbol{l}_{i}$ represents the channel estimation error of the corresponding estimation, $\omega_{i}$ is a set for all possible channel estimation errors, and $\Omega_{i}$ is the radii of the uncertainty regions as known to Alice. Hence, (2) can be reformulated as:
@@
$\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\triangleq\left(\widehat{{\boldsymbol{l}}}_{i}+\Delta\boldsymbol{l}_{i}\right)\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}}.$ (4)
@@
In the sequel, we deal with perfect and imperfect CSI from the IRS to the users and Eve. For the former, we assume that the CSI of all the users and Eve can be accurately obtained by Alice [23, 24]. As aforementioned, in this case we assume that Eve acts as a legitimate user in one period/slot but then tries to eavesdrop on other users. Using different well-established channel estimation methods, such as the anchor-assisted channel estimation approach [25], Alice can practically obtain the CSI of all the users. In this case, the reflected channel from the IRS to the users and Eve can be captured by setting $\Delta\boldsymbol{l}_{i}$ to zero in (3a), i.e., $\boldsymbol{l}_{i}=\widehat{\boldsymbol{l}}_{i},i\in\{k,e\}$. In the second case, only partial/imperfect CSI $\widehat{\boldsymbol{l}}_{i}$ is available. In this case, one can vary the magnitude of $\Delta\boldsymbol{l}_{i}$, i.e., the radii $\Omega_{i}$ of the uncertainty region to capture different levels of CSI imperfection.
@@
$\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq{|\boldsymbol{h}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{k}|^{2}}/{\rho_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (5)
$\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq{|\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{k}|^{2}}/{\widehat{\rho}_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (6)
where $\rho_{i}\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}|\boldsymbol{h}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i}$, and $\widehat{\rho}_{i}\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}|\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i}$. Under the presence of Eve, the closed-form expressions for the FBR-SR of the user-$k$ under perfect and imperfect IRS to users/Eve CSI are defined as, respectively, [26, Eq.115]:
$\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\left[C_{k}-C_{e}-\xi_{k}\sqrt{V_{k}}-\xi_{e}\sqrt{V_{e}}\right]^{+},$ (7)
$\widehat{\mathcal{S}}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\left[\widehat{C}_{k}-\widehat{C}_{e}-\xi_{k}\sqrt{\widehat{V}_{k}}-\xi_{e}\sqrt{\widehat{V}_{e}}\right]^{+},$ (8)
where $x]^{+}\triangleq\text{max}[0,x]$, while the user’s data rate, and the eavesdropping rate of Eve decoding $s_{k}$, are respectively given by,